OpenStudy (anonymous):

What is the slope of the line that passes through the points (2,7) and (-1,3) ?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

k gimme a sec..

OpenStudy (anonymous):



hartnn (hartnn):

yep, go on

hartnn (hartnn):

3-7 = -4 -1-2 = -3 so the slope will be ?

OpenStudy (anonymous):

it would be 

OpenStudy (anonymous):

no positive

hartnn (hartnn):

-4/-3 is just 4/3 :)

OpenStudy (anonymous):

because if there are 2 negatives, it would be postive

OpenStudy (anonymous):

ok :) thanks

hartnn (hartnn):

correct and welcome ^_^